# A Localized Orthogonal Decomposition method for heterogeneous mixed-dimensional problems

### v1

@@ 6. Practical multiscale method
We emphasize that the operator $\mathcal{R}^{\ell}$ depends on its argument only through its QOI, as introduced in 3.1. Consequently, the definition 6.1 of $\tilde{\varphi}^{\ell}_{T^{c}}$ is independent of the specific choice of bubble functions, provided the Kronecker-delta condition 3.6 on the QOI is satisfied. Note that, for a bulk element $T^{0}\in\mathcal{T}_{H}^{0}$, computing the associated basis function $\tilde{\varphi}_{T^{0}}^{\ell}$ requires solving local problems of the form 5.3 for all bulk elements in the first-order patch surrounding $T^{0}$. Similarly, for an interface element $T^{1}\in\mathcal{T}_{H}^{1}$, the basis function $\tilde{\varphi}_{T^{1}}^{\ell}$ is obtained by solving local problems of the form 5.3 for each bulk element that shares the face $T^{1}$. Figure 6.1 shows examples of a bulk basis function $\tilde{\varphi}_{T^{0}}^{\ell}$ and an interface basis function $\tilde{\varphi}_{T^{1}}^{\ell}$ for $\ell=4$.
@@
We emphasize that, for a practical computation of the LOD basis functions defined in 6.1, the problems 5.3, which are local but infinite-dimensional, still need to be solved. In practice, we approximate the solutions to these problems using a (local) fine-scale discretization. For codimensions $c\in\{0,1\}$, let $\mathcal{T}_{h}^{c}$ denote a mesh obtained by uniform refinement of the coarse mesh $\mathcal{T}_{H}^{c}$. The fine mesh size $h<H$ should be chosen such that $\mathcal{T}_{h}^{c}$ resolves the fine-scale oscillations of the coefficients of the problem at hand. Throughout the derivation of the method, we replace the space $V_{0}$ by its finite-dimensional subspace $V_{0,h}$, consisting of continuous, piecewise linear functions defined with respect to $\mathcal{T}_{h}^{c}$. Specifically, for problem 5.3, this means posing it on a finite-dimensional version of $V_{T^{0}}^{\ell}$ as defined in 5.1, but with $V_{0}$ replaced by $V_{0,h}$. In the fully discrete convergence analysis, $V_{0}$ is likewise replaced by $V_{0,h}$, and most arguments carry over directly; see, e.g., [31, Ch. 4.4], for a discussion of technical details that also arise in the present setting. This yields an a priori error estimate for the fully discrete LOD approximation similar to Theorem 6.1, but with respect to the fine-scale finite element solution in $V_{0,h}$. An estimate with respect to the weak solution of the original PDE then follows from the triangle inequality combined with standard finite element approximation results; see, e.g., [23, Thm. 2] for the present mixed-dimensional setting.
@@ 8. Numerical experiments
In this section, we present a numerical investigation of the proposed multiscale method. The first two numerical examples examine the convergence of the energy norm of the error, $\|u_{h}-\tilde{u}_{H}^{\ell}\|_{a}$, with respect to the coarse mesh size $H$ for the localized method defined in 6.2. Here, $u_{h}$ denotes the solution obtained from the fitted finite element method on the mesh $\mathcal{T}_{h}$, as introduced in [23]. The first example considers a problem with smoothly varying coefficients, while the second addresses rapidly oscillating coefficients $A^{0}$ and $A^{1}$. The third example investigates the localization error of the method, measured in the energy norm, with respect to the oversampling parameter $\ell$, for a coarse mesh containing complex-shaped elements. In this case, a simplified variant of method 6.2 is employed. Throughout this section, $H$ and $h$ denote the coarse and fine mesh sizes, respectively.
@@
Figure 8.1. The domain $\Omega$ used for the numerical experiments. Here, the blue lines are the interfaces $\Omega^{1}$ and the coarse mesh corresponding to $H=\tfrac{1}{4}$ is shown in red. The fine scale mesh is shown in the zoomed-in part.
@@
Figure 8.2. In 2a, the solution to the problem defined by $A^{0}=1$, $A^{1}=1$, $B^{1}=1$, $f^{0}=\sin(\pi x)\sin(\pi y)$, and $f^{1}=x+2y$ is shown. In 2b, we see the convergence results, the energy norm of the error between the reference solution $u_{h}$ and approximations $\tilde{u}_{H}^{\ell}$ with respect to $H$ is shown for different oversampling parameters $\ell$.
In the second numerical example, we consider coefficients $A^{0}$ and $A^{1}$ that vary rapidly on the fine scale. The coefficient $A^{0}$ is chosen to be piecewise constant on the fine mesh $\mathcal{T}_{h}$, taking random values in the range $[0.01,1]$, while $A^{1}$ is defined as $A^{1}=\sin(30\pi x)\sin(30\pi y)+2$. For this example, again we let $B^{1}=1$, $f^{0}\penalty 10000\ =\penalty 10000\ \sin(\pi x)\sin(\pi y)$, and $f^{1}=x+2y$. Figure 8.33a shows the corresponding finite element solution, and 3b presents the resulting convergence behavior. The observed convergence is qualitatively similar to that of the previous numerical example.
@@
Figure 8.3. In 3a, the solution of the problem with random, piecewise constant $A^{0}$ on the fine mesh, with random values in the range $0.01,\penalty 10000\ 1]$. Here, $A^{1}=\sin(30\pi x)\sin(30\pi y)+2$, $B^{1}=1$, $f^{0}=\sin(\pi x)\sin(\pi y)$, and $f^{1}=x+2y$. In 3b, the convergence of said problem with regards to the coarse mesh size $H$, for different oversampling parameters $\ell$ is shown.
In the final numerical example, we consider a generalized coarse mesh consisting of complex-shaped elements; see Section 7. The domain and corresponding coarse mesh are shown in Figure 7.11b. In this experiment, $A^{0}$ is again chosen to be piecewise constant on the fine mesh $\mathcal{T}_{h}$, with random values in the range $[0.01,1]$, while $A^{1}=1$, $B^{1}=1$, $f^{0}=1$, and $f^{1}=1$. The construction of the method for these complex-shaped coarse meshes follows the procedure described in Section 7. For implementation purposes, we deviate from the stabilized localization strategy presented in Section 5 and instead employ a simplified localization approach, where the global problems 3.5 defining the basis functions are localized directly. Note that this method is affected by the previously described deterioration of errors if the oversampling parameter is not increased logarithmically as the mesh size decreases. The resulting solution and convergence behavior for $h=\tfrac{1}{32}$ and $H=\tfrac{1}{2}$, with respect to the oversampling parameter $\ell$, are shown in Figure 8.4. We emphasize that for this choice of $f^{0}$ and $f^{1}$, the first term on the right-hand side of error estimate 6.3 vanishes and only the exponentially decaying localization error remains. The observed decay behavior is again consistent with Theorem 3.2.

### v2

@@ 6. Practical multiscale method
We emphasize that the operator $\mathcal{R}^{\ell}$ depends on its argument only through its QOI, as introduced in 3.1. Consequently, the definition 6.1 of $\tilde{\varphi}^{\ell}_{T^{c}}$ is independent of the specific choice of bubble functions, provided that the Kronecker-delta condition 3.6 on the QOI is satisfied. Note that, for a bulk element $T^{0}\in\mathcal{T}_{H}^{0}$, computing the associated basis function $\tilde{\varphi}_{T^{0}}^{\ell}$ requires solving local problems of the form 5.3 for all bulk elements in the first-order patch surrounding $T^{0}$. Similarly, for an interface element $T^{1}\in\mathcal{T}_{H}^{1}$, the basis function $\tilde{\varphi}_{T^{1}}^{\ell}$ is obtained by solving local problems of the form 5.3 for each bulk element that shares the face $T^{1}$. Figure 6.1 shows examples of a bulk basis function $\tilde{\varphi}_{T^{0}}^{\ell}$ and an interface basis function $\tilde{\varphi}_{T^{1}}^{\ell}$ for $\ell=4$.
@@
We emphasize that, for a practical computation of the LOD basis functions defined in 6.1, the problems 5.3, which are local but infinite-dimensional, still need to be solved. In practice, we approximate the solutions to these problems using a (local) fine-scale discretization. For codimensions $c\in\{0,1\}$, let $\mathcal{T}_{h}^{c}$ denote a mesh obtained by uniform refinement of the coarse mesh $\mathcal{T}_{H}^{c}$. The fine mesh size $h<H$ should be chosen such that $\mathcal{T}_{h}^{c}$ resolves the fine-scale oscillations of the coefficients of the problem at hand. Throughout the derivation of the method, we replace the space $V_{0}$ by its finite-dimensional subspace $V_{0,h}$, consisting of continuous, piecewise linear functions defined with respect to $\mathcal{T}_{h}^{c}$. Specifically, for problem 5.3, this means posing it on a finite-dimensional version of $V_{T^{0}}^{\ell}$ as defined in 5.1, but with $V_{0}$ replaced by $V_{0,h}$. In the fully discrete convergence analysis, $V_{0}$ is likewise replaced by $V_{0,h}$, and most arguments carry over directly; see, e.g., [31, Ch. 4.4], for a discussion of technical details that also arise in the present setting. This yields an a priori error estimate for the fully discrete LOD approximation similar to theorem 6.1, but with respect to the fine-scale finite element solution in $V_{0,h}$. An estimate with respect to the weak solution of the original PDE then follows from the triangle inequality combined with standard finite element approximation results; see, e.g., [25, Thm. 2] for the present mixed-dimensional setting.
@@ 8. Numerical experiments
In this section, we present a numerical investigation of the proposed multiscale method. The first two numerical examples examine the convergence of the energy norm of the error, $\|u_{h}-\tilde{u}_{H,h}^{\ell}\|_{a}$, with respect to the coarse mesh size $H$ for the localized method defined in 6.2. Here, $u_{h}$ denotes the solution obtained from the fitted finite element method on the mesh $\mathcal{T}_{h}$, as introduced in [25], and $\tilde{u}_{H,h}^{\ell}$ denotes the solution to the discretized version of problem 6.2. The first example considers a problem with smoothly varying coefficients, while the second addresses rapidly oscillating coefficients $A^{0}$ and $A^{1}$. The third example investigates the localization error of the method, measured in the energy norm, with respect to the oversampling parameter $\ell$, for a coarse mesh containing complex-shaped elements. In this case, a simplified variant of method 6.2 is employed. Finally, in the last example we use a more complex domain and an example that highlights the multiscale features more clearly. In this example, we look at the convergence following $\ell=\log_{2}(1/H)$ in order to numerically verify the theoretical results from theorem 6.1. Throughout this section, $H$ and $h$ denote the coarse and fine mesh sizes, respectively.
@@
Figure 8.1. The domain $\Omega$ used for the first two numerical experiments. Here, the blue lines are the interfaces $\Omega^{1}$ and the coarse mesh corresponding to $H=\tfrac{1}{4}$ is shown in red. The fine-scale mesh is shown in the zoomed-in part.
@@
Figure 8.2. In 2a, the solution to the problem defined by $A^{0}=1$, $A^{1}=1$, $B^{1}=1$, $f^{0}=\sin(\pi x)\sin(\pi y)$, and $f^{1}=x+2y$ is shown. In 2b, we see the convergence results, the energy norm of the error between the reference solution $u_{h}$ and approximations $\tilde{u}_{H,h}^{\ell}$ with respect to $H$ is shown for different oversampling parameters $\ell$.
In the second numerical example, we consider coefficients $A^{0}$ and $A^{1}$ that vary rapidly on the fine scale. The coefficient $A^{0}$ is chosen to be piecewise constant on the fine mesh $\mathcal{T}_{h}$, taking random values in the range $[0.01,1]$, while $A^{1}$ is defined as $A^{1}=\penalty 10000\ \sin(30\pi x)\sin(30\pi y)+2$. For this example, again we let $B^{1}=1$, $f^{0}\penalty 10000\ =\penalty 10000\ \sin(\pi x)\sin(\pi y)$, and $f^{1}=x+2y$. figure 8.33a shows the corresponding finite element solution, and 3b presents the resulting convergence behavior. The observed convergence is qualitatively similar to that of the previous numerical example.
@@
Figure 8.3. In 3a, the solution of the problem with random, piecewise constant $A^{0}$ on the fine mesh, with random values in the range $0.01,\penalty 10000\ 1]$. Here, $A^{1}=\penalty 10000\ \sin(30\pi x)\sin(30\pi y)+2$, $B^{1}=1$, $f^{0}=\sin(\pi x)\sin(\pi y)$, and $f^{1}=x+2y$. In 3b, the convergence of said problem with regards to the coarse mesh size $H$, for different oversampling parameters $\ell$ is shown.
In the third numerical example, we consider a generalized coarse mesh consisting of complex-shaped elements; see section 7. The domain and corresponding coarse mesh are shown in figure 7.11b. In this experiment, $A^{0}$ is again chosen to be piecewise constant on the fine mesh $\mathcal{T}_{h}$, with random values in the range $[0.01,1]$, while $A^{1}=1$, $B^{1}=1$, $f^{0}=1$, and $f^{1}=1$. The construction of the method for these complex-shaped coarse meshes follows the procedure described in section 7. For implementation purposes, we deviate from the stabilized localization strategy presented in section 5 and instead employ a simplified localization approach, where the global problems 3.5 defining the basis functions are localized directly. Note that this method is affected by the previously described deterioration of errors if the oversampling parameter is not increased logarithmically as the mesh size decreases. The resulting solution and convergence behavior for $h=\tfrac{1}{32}$ and $H=\tfrac{1}{2}$, with respect to the oversampling parameter $\ell$, are shown in figure 8.4. We emphasize that for this choice of $f^{0}$ and $f^{1}$, the first term on the right-hand side of error estimate 6.3 vanishes and only the exponentially decaying localization error remains. The observed decay behavior is again consistent with theorem 3.2.
@@
In the final numerical example, we instead use the domain shown in Figure 8.5, utilising more interfaces for higher complexity. We let $A_{0}$ be piecewise constant on the fine mesh with random values in $0.01,\penalty 10000\ 1]$, $A^{1}=\sin(30\pi x)\sin(30\pi y)\penalty 10000\ +\penalty 10000\ 1.1$, $B^{1}=1$, $f^{0}=\sin(\pi x)\sin(\pi y)$ and $f^{1}=x+2y$. Altering $A^{1}$ in this manner, we can more clearly see the multiscale effects of the solution on the interfaces, see figure 8.6 6a. figure 8.6 6b shows the convergence with regards to $H$, letting $\ell=\log_{2}(1/H)$. As expected from theorem 6.1, we obtain $H^{2}$-convergence using the smooth right hand side. For contrast, we also let $f^{0,1}=\sin(30\pi x)\sin(\pi y)$, which is rapidly varying and not resolved by $H$ but is resolved by $h$. For this case, we get a lower order of convergence, which is expected according to theorem 6.1.
Figure 8.5. The domain $\Omega$ used for the final numerical experiment. Here, the blue lines are the interfaces $\Omega^{1}$ and the coarse mesh corresponding to $H=\tfrac{1}{4}$ is shown in red.
a
[…]
b
Figure 8.6. In 6a, the finite element solution is shown that corresponds to the problem with piecewise constant $A^{0}$ on the fine mesh, with values in the range $0.01,\penalty 10000\ 1]$. Here, $A^{1}=\sin(30\pi x)\sin(30\pi y)+1.1$ in order for the multiscale features to appear more clearly, whereas the other data is as in the previous examples. In 6b, we see the convergence with respect to $H$, following $\ell=\log_{2}(\frac{1}{H})$. The convergence for a smooth right hand side follows $H^{2}$ whereas using a more rapidly varying right hand side, we obtain a lower order of convergence.